OpenStudy (anonymous):

If x represents the probability of an event happening, then which of the following represents the probability that the event will not happen? -x 1 - x x - 1

jimthompson5910 (jim_thompson5910):

which one do you think it is?

OpenStudy (anonymous):

-x

jimthompson5910 (jim_thompson5910):

say x is 30% so -x is -30% is that possible?

jimthompson5910 (jim_thompson5910):

a negative probability is possible?

jimthompson5910 (jim_thompson5910):

are you sure?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

it turns out that probabilities must be positive

OpenStudy (anonymous):

it would make it the bottom one to come out positive

jimthompson5910 (jim_thompson5910):

are you referring to x-1?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

well if x = 0.3, then x -1 = -0.7, which is also not possible

OpenStudy (anonymous):

oh now i see the second

jimthompson5910 (jim_thompson5910):

yes, ex: if x = 0.4, then 1-x = 1-0.4 = 0.6